# Intensity-Efficient Allocations

### v5

@@ 5 Eliciting Ordinal Intensities and Finding Intensity-Dominating Allocations
Define $f(n):=\frac{1}{2}(n-1)(n-2)$. Observe that (4) and $(\mathbf{5}_{D})$ together imply that any $\dot{\succsim}_{l}\in\mathcal$$\widehat${I}} contains
$|\dot{\succsim}_{l}|$ $=$ $2{n\choose 2}+n$ (31)
comparisons between pairs in $A\times A$. We proceed by induction.
Let $n=3$ and, without loss, $a_{1}\succ_{l}a_{2}\succ_{l}a_{3}$. By (31), $|\dot{\succsim}_{l}|=9$. By (6) and the assumption that $\succ_{l}$ is induced by $\dot{\succsim}_{l}$ we readily get $(a_{1},a_{3})\dot{\succsim}_{l}(a_{1},a_{2})$ and $(a_{1},a_{3})\dot{\succsim}_{l}(a_{2},a_{3})$. By (5), moreover, we have $(a_{1},a_{1})\dot{\sim}_{l}(a_{2},a_{2})\dot{\sim}_{l}(a_{3},a_{3})$. By (5), $(\mathbf{5}_{D})$ and $a_{1}\succ_{l}a_{2}\succ_{l}a_{3}$ we also get $(a_{1},a_{2})\dot{\succ}_{l}(a_{i},a_{i})$ and $(a_{2},a_{3})\dot{\succ}_{l}(a_{i},a_{i})$, $i\leq 3$. By (5) and the above comparisons, finally, we also have $(a_{i},a_{i})\dot{\succ}_{l}(a_{2},a_{1})\dot{\succ}_{l}(a_{3},a_{1})$ and $(a_{i},a_{i})\dot{\succ}_{l}(a_{3},a_{2})\dot{\succ}_{l}(a_{3},a_{1})$. Together, these account for 7 out of the 9 comparisons in $\dot{\succsim}_{l}$, with those between $(a_{1},a_{2})$, $(a_{2},a_{3})$ and between $(a_{3},a_{2})$, $(a_{2},a_{1})$ being the two missing ones. By (5) once again, knowing either the first or the second such comparison also pins down the other. Hence, $f(3)=1$ is indeed the number of steps needed to elicit $\dot{\succsim}_{l}$ from $\succ_{l}$ when $n=3$.
Now let $n=4$ and, without loss, $a_{1}\succ_{l}a_{2}\succ_{l}a_{3}\succ_{l}a_{4}$. By the preceding arguments we know that, out of the $|\dot{\succsim}_{l}|=16$ total comparisons here, missing are the following 3: $(a_{1},a_{2})$ vs $(a_{2},a_{3})$; $(a_{2},a_{3})$ vs $(a_{3},a_{4})$; $(a_{1},a_{3})$ vs $(a_{2},a_{4})$. Hence, $f(4)=3$ is the number of necessary steps here too.
For the induction step of the proof when $n>4$, suppose $f(m)$ is the correct number for every $n\leq m$. We will show that $f(m+1)$ is the correct number when $n=m+1$. To this end, let $a_{1}\succ_{l}a_{2}\succ_{l}\ldots\succ_{l}a_{m}$ and notice that, from the induction hypothesis and the information derived from the $n=3$ and $n=4$ cases above, we can decompose the $f(m)$ comparisons as follows:
(1):  $(a_{i},a_{i+1})$ vs $(a_{i+1},a_{i+2})$ for $i\leq m-2$ (m2m-2 comparisons);
(2):  $(a_{i},a_{i+2})$ vs $(a_{i+1},a_{i+3})$ for $i\leq m-3$ (m3m-3 comparisons);
(3):  $(a_{i},a_{i+3})$ vs $(a_{i+1},a_{i+4})$ for $i\leq m-4$ (m4m-4 comparisons);
$\cdots$
(m2m-2):  $(a_{i},a_{i+m-1})$ vs $(a_{i+1},a_{i+m})$ (1 comparison).
Indeed, summing the terms inside brackets across the right hand sides of $(1)\cdots(m-2)$ we get
$\sum\limits_{i=2}^{m-1}(m-i)$ $=$ $\sum\limits_{i=1}^{m-1}(m-i)-(m-1)$
$=$ $\frac{1}{2}(m-1)m-(m-1)$
$=$ $(m-1)\left(\frac{1}{2}m-1\right)$
$=$ $\frac{1}{2}(m-1)(m-2)$
$=$ $f(m)$
Consider now $n=m+1$ and, once again without loss, let $a_{1}\succ_{l}a_{2}\succ_{l}\ldots\succ_{l}a_{m+1}$. From the preceding decomposition of the required comparisons when $n=m$ we can deduce that, for $n=m+1$,
(1):  $(a_{i},a_{i+1})$ vs $(a_{i+1},a_{i+2})$ for $i\leq m-1$ (m1m-1 comparisons);
(2):  $(a_{i},a_{i+2})$ vs $(a_{i+1},a_{i+3})$ for $i\leq m-2$ (m2m-2 comparisons);
(3):  $(a_{i},a_{i+3})$ vs $(a_{i+1},a_{i+4})$ for $i\leq m-3$ (m3m-3 comparisons);
$\cdots$
(m1m-1):  $(a_{i},a_{i+m})$ vs $(a_{i+1},a_{i+m+1})$ (1 comparison).
Summing across categories now yields $\sum\limits_{i=1}^{m}(m-i)=\frac{1}{2}(m-1)m=f(m+1)$. ∎
Turning now to the problem of using as input a strict intensity profile SS—and the strict preference profile $P_{S}$ nested within it—to find an allocation that improves upon Pareto-efficiency, we proceed by recalling first some important landmarks from the matching-theoretic literature with one-sided preferences where no intensity information is available. The study of such rules goes back to Shapley and Scarf (1974) and David Gale’s Top-Trading Cycles (TTC) algorithm. The premise there is that each agent has an initial endowment from the set of items that will be (re-)assigned once the algorithm is run. TTC then operates sequentially by implementing swaps between these initial assignments in accordance with the agents’ preferences, giving priority to an item’s current assignee in situations where additional agents prefer this over their own endowment. Shapley and Scarf (1974) proved that TTC leads to a core allocation. Roth and Postlewaite (1977) later showed that this core allocation, identifiable by TTC, is unique. Extending the results in Roth (1982), Ma (1994) later established that TTC is the unique rule that leads to a Pareto-efficient allocation and, in addition, satisfies strategy-proofness and individual-rationality.
Hylland and Zeckhauser (1979) introduced randomness to study the assignment problem without initial endowments and, assuming that all agents have an equal amount of some artificial currency and expected-utility preferences over lotteries over allocations, studied that which is associated with (or approximated by) Competitive Equilibrium from Equal Incomes. This mechanism is not strategy-proof. Also introducing randomness, but without assuming expected-utility preferences or employing the competitive-equilibrium apparatus, Abdulkadiroğlu and Sönmez (1998) showed that TTC with random initial endowments is equivalent to the Random Priority (RP) mechanism. The latter does not involve initial endowments; ranks agents under a random priority order; and lets them choose, accordingly, their most preferred item from those still available. These authors also clarified that a deterministic matching mechanism leads to a Pareto-efficient allocation if and only if is a Random Priority under some ranking over agents.
Bogomolnaia and Moulin (2001), finally, demonstrated that, despite the ex post Pareto-efficiency of RP allocations, hence of core/TTC allocations under random endowments, these are generally ex ante inefficient in the sense that, based on the stochastic-dominance relation defined by all agents’ preferences, they may be dominated by another random allocation. In response to this realization, the authors introduced ordinally-efficient allocations as those that are not first-order stochastically dominated by any other random allocation, and proved that their Probabilistic Serial (PS) mechanism, although not strategy-proof, leads to an ordinally-efficient allocation with desirable—under expected-utility preferences over allocations—ex ante fairness properties.
Input: A strict intensity profile $S\in\mathcal{\widehat{S}}$.
Step 1: Choose a linear order $\gg$ over $V$ from the uniform distribution with support $n!$.
Step 4: Search for the $\gg$-lowest agent $v_{j}\in\{v_{2},\ldots,v_{n}\}$ such that $s_{1}(x_{j},x_{1})>s_{j}(x_{j},x_{1})$. If such an agent exists, swap $x_{1}\leftrightarrow x_{j}$ and remove $v_{1},v_{j}$ from $V$ and $x_{1},x_{j}$ from $A$.
If no such agent exists, remove $v_{1}$, $x_{1}$.
Repeat Step 4 until no agents remain.
Proposition 4.
If $S\in\mathcal{\widehat{S}}$ and either $n=3$ or $S$ is effectively homeo-monotonic, then the algorithm terminates in $O(n^{2})$ time and either produces the Pareto-efficient allocation that it finds in its RP-implementing Step 2 or an allocation that intensity-dominates it.
@@
Let $x$ be the Pareto-efficient allocation that is found in Step 2 under the random priority ordering $\gg$ drawn in the algorithm’s first step. In the worst-case scenario, Step 4 always—with the possible exception of the agent ranked second-last by $\gg$—performs exhaustive search and removes one agent and one object in every iteration: Starting with $v_{1},x_{1}$, after n1n-1 searches it removes $v_{1},x_{1}$. Continuing with $v_{2},x_{2}$, after n2n-2 searches it removes $v_{2},x_{2}$. $\cdots$ Continuing with $v_{n-1},x_{n-1}$, after one search it removes both $v_{n-1},x_{n-1}$ and $v_{n},x_{n}$, either by implementing the swap $x_{n-1}\leftrightarrow x_{n}$ or otherwise. Worst-case performance is therefore associated with a total number of searches equalling
$\sum_{m=1}^{n-1}(n-m)$ $=$ $n^{2}-n-\sum_{m=1}^{n-1}m$
$=$ $n^{2}-\sum_{m=1}^{n}m$
@@
Given that a Pareto-efficient allocation can be found by RP in Step 2 in $O(n)$ time (Manlove, 2013, Section 6.2), it follows that the algorithm terminates in $O(n^{2})$ time.
@@
It remains to be shown that, if the final allocation is $y\neq x$, then $yDx$. Suppose $n=3$ first. Write $x:=(x_{1},x_{2},x_{3})$. By the definition of Step 4, the transition from $x$ to $y$ involves a single pairwise swap. Thus, $y$ must be one of the following:
(i) $(x_{1},x_{3},x_{2})$, implying $s_{2}(x_{3},x_{2})>s_{3}(x_{3},x_{2})>0$ or $s_{3}(x_{2},x_{3})>s_{2}(x_{2},x_{3})>0$
(ii) $(x_{3},x_{2},x_{1})$, implying $s_{1}(x_{3},x_{1})>s_{3}(x_{3},x_{1})>0$ or $s_{3}(x_{1},x_{3})>s_{1}(x_{1},x_{3})>0$
(iii) $(x_{2},x_{1},x_{3})$, implying $s_{2}(x_{1},x_{2})>s_{1}(x_{1},x_{2})>0$ or $s_{1}(x_{2},x_{1})>s_{2}(x_{2},x_{1})>0$
@@
In each case, $yDx$. Now suppose $n>3$ and $S\in\mathcal{\widehat{S}}$ is effectively homeo-monotonic. Denote by $(V^{i})_{i=1}^{k}$ and $(A^{i})_{i=1}^{k}$ the respective partitions of $V$ and $A$ into a (possibly empty) top-preference diverse set $V^{1}$ with respect to $A^{1}$ and top-preference-coincident sets $V^{i>1}$ with respect to $A^{i>1}$, $i\leq k$. By the properties of these sets profiles, and by the definition of algorithm’s Step 4, irrespective of the $\gg$-order on $V$ there can be no swaps between $x_{i}$ and $x_{j}$ for $v_{i},v_{j}\in V^{1}$. Thus, $y_{i}=x_{i}$ for all $v_{i}\in V^{1}$. By the properties of $(V^{i>1},A^{i>1})$, moreover and the instruction in Step 4, there can be no swaps between $x_{i}$ and $x_{j}$ for $v_{i}\in V^{l>1}$ and $v_{j}\in V^{m>1}$ when $m,l\in\{2,\ldots,k\}$ are distinct. It follows that $y$ differs from $x$ by at most $\frac{n}{2}$ or $\frac{n-1}{2}$ pairwise swaps/transpositions if $n$ is even or odd, respectively, with each such swap $x_{i}\leftrightarrow x_{j}$ occurring within the same top-preference-coincident pair $(V^{l>1},A^{l>1})$. By Step 4, finally, every such swap reassigns the jointly preferred (by Pareto-efficiency of $x$) item within the relevant pair in the direction of intensity-dominance. Hence, $yDx$. ∎
The allocation delivered by this algorithm is not intensity-efficient in general. This is so because the forced exit of agents and items in the fourth step—a crucial feature for the algorithm’s manageable quadratic-time complexity—means that not all pairwise possibilities for intensity-dominance-driven improvements are considered in the transition from the RP-mandated Pareto-efficient allocation in Step 2 to the final allocation. Clearly, however, the algorithm improves upon RP and Pareto-efficiency—even if partially so—in the direction of intensity-efficiency, without adding intractable computational complexity relative to RP. It is presently unknown if this algorithm can be modified to deliver an intensity-efficient allocation—whenever one is known to exist—without requiring factorial search over all possible priority orderings over agents in Step 2.
In light of the discussion preceding the statement of Proposition 4, it is clear that RP can be replaced by PS in Step 2, depending on the desired resolution between strategy-proofness and ordinal-efficiency. Regardless of whether RP, PS or some other mechanism is used to elicit the preference component of the input intensity profile in the manner dictated in the proof of Proposition 3, however, one is naturally interested to know if, when operating under the matching rule defined by the above algorithm, agents are incentivized to report their intensities truthfully. The answer is negative. To see this, suppose $n=3$ and let $S\in\mathcal{\widehat{S}}$ be such that $a\succ_{1}b\succ_{1}c$; $a\succ_{2}b\succ_{2}c$; $c\succ_{3}b\succ_{3}a$; $(a,b)\dot{\succ}_{1}(b,c)$; $(b,c)\dot{\succ}_{2}(a,b)$; $(c,b)\dot{\succ}_{3}(b,a)$. The possible orderings over $V$ are: 1. $v_{1}\gg v_{2}\gg v_{3}$; 2. $v_{1}\gg v_{3}\gg v_{2}$; 3. $v_{2}\gg v_{1}\gg v_{3}$; 4. $v_{2}\gg v_{3}\gg v_{1}$; 5. $v_{3}\gg v_{1}\gg v_{2}$; 6. $v_{3}\gg v_{2}\gg v_{1}$. The two Pareto-efficient allocations are $(a,b,c)$ and $(b,a,c)$, and the former intensity-dominates the latter. Suppose the algorithm’s second step delivers this intensity-efficient allocation (under RP, this happens with probability $\frac{1}{2}$). If intensities are reported truthfully, there is no order $\gg$ under which the algorithm’s subsequent steps alter it. However, if $(b,c)\dot{\succ}^{\prime}_{2}(a,b)$ is the second agent’s misreported intensity, then under the 3rd, 4th and 6th $\gg$-orders (i.e. with probability $\frac{1}{2}$) the algorithm reassigns $b$ to $v_{1}$ and $a$ to $v_{2}$, making the latter agent better off.
Interestingly, the incentives to misreport one’s intensities in this way, namely toward increasing one’s chances of getting their most preferred alternative, act in the opposite direction to the incentives toward misreporting one’s preferences in the Boston/Immediate-Acceptance mechanism of the distinct two-sided (School Choice) matching problem. Specifically, as has been discussed by Pathak and Sönmez (2008); Abdulkadiroğlu, Che, and Yasuda (2011) and other authors, under this mechanism students may (and often do) submit their second- or third-ranked school as their supposedly most preferred one if their actual top-ranked school is very popular, out of concern that in the event they are not assigned to it they may end up at one of their lower-ranked schools if their second or third most preferred ones have in the meantime filled up. Abdulkadiroğlu et al. (2011), however, followed by Pycia (2011); Featherstone and Niederle (2016) and other studies, argued that this drawback on incentive grounds of the Boston mechanism may be outweighed by the total welfare gains associated with such misreporting once the students’ intensities are accounted for. In line with the prior relevant literature, intensity comparisons in that paper’s Bayesian-Nash equilibrium analysis are derived from intra- and inter-personal differences in cardinal utility functions that emerge from expected-utility preferences over lotteries over schools. In that framework too, finally, the intensity information that is conveyed to the social planner is assumed to exist without having been elicited incentive-compatibly, and it is unclear if such elicitation is possible. In relation to this, Pycia and Ünver (2025) recently showed that, in environments without monetary transfers such as the one considered above, only ordinal mechanisms are group strategy-proof.

### v6

@@ 6 Eliciting Ordinal Intensities and Improving Upon Pareto-Efficient Allocations
By definition,
$\begin{array}[]{ccccc}r_{l}(a_{i})\equiv\max\limits_{a_{k}\in A}s_{l}(a_{i},a_{k})&\geq&r_{l}(a_{j})\equiv\max\limits_{a_{k}\in A}s_{l}(a_{j},a_{k})&\Longleftrightarrow\\ s_{l}(a_{i},a_{k})&\geq&s_{l}(a_{j},a_{k})\quad\forall\;a_{k}\in A&\Longleftrightarrow\\ (a_{i},a_{k})&\dot{\succsim}_{l}&(a_{j},a_{k})\quad\forall\;a_{k}\in A&\Longleftrightarrow&\\ (a_{i},a_{j})&\dot{\succsim}_{l}&(a_{j},a_{i})&\Longleftrightarrow&a_{i}\succsim_{l}a_{j},\end{array}$
where the penultimate equivalence immediately follows from $(\mathbf{4_{\mathit{D}}})$, which in turn is implied by the postulate that $s_{l}$ represents $\dot{\succsim}_{l}$ in profile $S\in\mathcal{\widetilde{S}}$. ∎
Considering now the stronger postulate whereby $\dot{\succsim}_{l}\in\mathcal$$\widetilde${I}} is linear, note first that this implies that the relation $\succsim_{l}$ induced by $\dot{\succsim}_{l}$ is a linear order. To ease notation, in the rest of the proof we assume without loss that the $n$ elements of $A$ are $\succ_{l}$-ranked by every agent $l\in N$ as follows:
$\begin{array}[]{lllllllllll}a_{1}&\succ_{l}&a_{2}&\succ_{l}&a_{3}&\succ_{l}&\cdots&\succ_{l}&a_{n-1}&\succ_{l}&a_{n}\end{array}$
(Dealing with heterogeneous preferences in $P_{S}$ is straightforward and involves replacing the index $i$ in $a_{i}$ with $\pi^{l}(i)$ for a suitable agent-specific permutation $\pi^{l}$ on $\{1,\ldots,n\}$.)
Lemma 6.
[…]
$r_{l}(A)$ $=$ $\{0,1,\ldots,n-1\}$ (38)
@@
Consider a linear $\dot{\succsim}_{l}\in\mathcal$$\widetilde${I}} and assume that it induces the linear order $\succ_{l}$ above. We will show that the “upper half” of this intensity ordering—in which the left item in each pair is preferred to its right counterpart—exhibits the following pattern:
$(a_{1},a_{n})$
$\dot{\succ}_{l}$
$(a_{1},a_{n-1})\,\dot{\sim}_{l}\,(a_{2},a_{n})$
$\vdots$
$(a_{1},a_{3})\,\dot{\sim}_{l}\,(a_{2},a_{4})\,\dot{\sim}_{l}\,\cdots\,\dot{\sim}_{l}\,(a_{n-2},a_{n})$
$\dot{\succ}_{l}$
$(a_{1},a_{2})\,\dot{\sim}_{l}\,(a_{2},a_{3})\,\dot{\sim}_{l}\,(a_{3},a_{4})\,\dot{\sim}_{l}\,\cdots\,\dot{\sim}_{l}\,(a_{n-1},a_{n})$
$\dot{\succ}_{l}$
$(a_{1},a_{1})\,\dot{\sim}_{l}\,(a_{2},a_{2})\,\dot{\sim}_{l}\,(a_{3},a_{3})\,\dot{\sim}_{l}\,\cdots\,\dot{\sim}_{l}\,(a_{n-1},a_{n-1})\,\dot{\sim}_{l}\,(a_{n},a_{n})$
The bottom two lines are true by reflexivity of $\dot{\succsim}_{l}$ and by the postulated linearity of $\dot{\succsim}_{l}$, respectively, while the $\dot{\succ}_{l}$ sign between them is implied by $a_{i}\succ_{l}a_{i+1}$ for all $i\leq n-1$. Consider now the preceding line. We know from the above that $s_{l}(a_{i},a_{i+1})=s_{l}(a_{i+1},a_{i+2})=s_{l}(a_{i+2},a_{i+3})>0$. Hence, by (6), $s_{l}(a_{i},a_{i+2})=s_{l}(a_{i+1},a_{i+2})$ is also true. Since $s_{l}(a_{i+1},a_{i+2})=s_{l}(a_{i+2},a_{i+3})$ holds too, (6) further implies $s_{l}(a_{i+1},a_{i+3})=s_{l}(a_{i+1},a_{i+2})$. Combined, these yield $s_{l}(a_{i},a_{i+2})=s_{l}(a_{i+1},a_{i+3})$ or, equivalently, $(a_{i},a_{i+2})\dot{\sim}_{l}$ $(a_{i+1},a_{i+3})$. This proves that the intensity-equivalence statements in that line are valid too, while the $\dot{\succ}_{l}$ sign following it is a direct consequence of (6). By finiteness, proceeding upwards in this fashion and applying the same argument repeatedly establishes the claimed pattern in the structure of $\dot{\succsim}_{l}$. (Equivalently, this pyramidal scheme shows how the top-half intensity-equivalence classes of $\dot{\succsim}_{l}$ are formed and ordered by $\dot{\succ}_{l}$.)
Since there are $n$ alternatives in $A$, the above shows that there are n1n-1 intensity-equivalence classes $a_{i},a_{j}]\in(A\times A)_{\dot{\sim}_{l}}$ such that $a_{i}\succ_{l}a_{j}$. By ($1$), $r_{l}(a_{i})\equiv\max\limits_{a_{k}\in A}s_{l}(a_{i},a_{k})$. This and the above-postulated $\succ_{l}$ ordering on AA jointly imply
$s_{l}(a_{1},a_{n})$ $=$ $n-1$
$s_{l}(a_{2},a_{n})=s_{l}(a_{1},a_{n-1})$ $=$ $n-2$
$\vdots$
$s_{l}(a_{1},a_{2})=\cdots=s_{l}(a_{n-1},a_{n})$ $=$ $1$
$s_{l}(a_{1},a_{1})=\cdots=s_{l}(a_{n},a_{n})$ $=$ $0$
Thus, in line with Lemma 5, we have $a_{i}\succ_{l}a_{j}\Leftrightarrow r_{l}(a_{i})>r_{l}(a_{j})$ and, in addition,
[…]
from which the claim follows. ∎
Now consider the induced profile $P_{S}\in\mathcal{P}$. By definition, $b_{l}^{P_{S}}(a_{i})\equiv|\{a_{k}\in A:a_{i}\succ_{l}a_{k}\}|$. This and the given ordering $\succ_{l}$ on $A$ readily imply $b_{l}^{P_{S}}(a_{i})=n-i$. Therefore, for every linear intensity profile $S\in\mathcal{\widetilde{S}}$, every agent $l\leq n$, and every alternative $a_{i}\in A$,
$r_{l}^{S}(a_{i})$ $=$ $b_{l}^{P_{S}}(a_{i})$
$=$ $n-i$
[…]
Proof of Proposition 3. By the definitions of $b_{l}$ and $r_{l}$ in (10) and (11), for any $S\in\mathcal{\widehat{S}}$ and every $l\leq n$ there is a strictly increasing function $f_{l}:\{0,1,2,\ldots,n-1\}\rightarrow\{0,1,2,\ldots,{{n}\choose{2}}\}$ such that, for all $a\in A$,
$r_{l}^{S}(a)$ $\equiv$ $f_{l}(b_{l}^{P_{S}}(a))$
Hence, for $x,y\in\mathcal{A}$ it holds that
$\sum_{i=1}^{n}b_{l}^{P_{S}}(x_{i})>\sum_{i=1}^{n}b_{l}^{P_{S}}(y_{i})$ $\Longrightarrow$ $\sum_{i=1}^{n}r_{l}^{S}(x_{i})>\sum_{i=1}^{n}r_{l}^{S}(y_{i})$ (39)
To prove the first claim, suppose to the contrary that $x\in\mathsf{IP}(S)$ but $x\not\in\mathsf{Bo}(S)$. Then, there must be $y\in\mathcal{A}$ such that $\sum\limits_{i=1}^{n}b_{l}^{P_{S}}(y_{i})>\sum\limits_{i=1}^{n}b_{l}^{P_{S}}(x_{i})$. By (39), this contradicts the postulate $x\in\mathsf{IP}(S)$. The second claim is proved in the main text. $\square$
Proof of Proposition 4. In the main text.
Proof of Proposition 5. It suffices to note that (4) and $(\mathbf{5}_{D})$ together imply that any $\dot{\succsim}_{l}\in\mathcal$$\widehat${I}} contains $|\dot{\succsim}_{l}|=2{n\choose 2}+n=(n-1)(n+1)$ comparisons between pairs in $A\times A$. $\square$
@@
Proof of Proposition 6. Let $x$ be the Pareto-efficient allocation that is found in Step 2 under the random priority ordering $\gg$ drawn in the algorithm’s first step. In the worst-case scenario, Step 4 always—with the possible exception of the agent ranked second-last by $\gg$—performs exhaustive search and removes one agent and one object in every iteration: Starting with $v_{1},x_{1}$, after n1n-1 searches it removes $v_{1},x_{1}$. Continuing with $v_{2},x_{2}$, after n2n-2 searches it removes $v_{2},x_{2}$. $\cdots$ Continuing with $v_{n-1},x_{n-1}$, after one search it removes both $v_{n-1},x_{n-1}$ and $v_{n},x_{n}$, either by implementing the swap $x_{n-1}\leftrightarrow x_{n}$ or otherwise. Worst-case performance is therefore associated with a total number of searches equalling
$\begin{array}[]{llclcllll}\sum\limits_{m=1}^{n-1}(n-m)&=&n^{2}-n-\sum\limits_{m=1}^{n-1}m&=&n^{2}-\sum\limits_{m=1}^{n}m&=&&&\\ &&&&&&\\ &&n^{2}-\dfrac{n}{2}(n+1)&=&\dfrac{1}{2}n(n-1)&=&\displaystyle{{n\choose k}}&&\end{array}$
@@
Given that a Pareto-efficient allocation can be found by RP in Step 2 in $O(n)$ time (Manlove, 2013, Section 6.2), it follows that the algorithm terminates in at most ${n}\choose{2}$} steps, hence in $O(n^{2})$ time.
@@
It remains to be shown that, if the final allocation is $y\neq x$, then $yDx$. Suppose $n=3$ first. Write $x:=(x_{1},x_{2},x_{3})$. By the definition of Step 4, the transition from $x$ to $y$ involves a single pairwise swap. Thus, $y$ must be one of the following: (i) $(x_{1},x_{3},x_{2})$, implying $s_{2}(x_{3},x_{2})>s_{3}(x_{3},x_{2})>0$ or $s_{3}(x_{2},x_{3})>s_{2}(x_{2},x_{3})>0$; (ii) $(x_{3},x_{2},x_{1})$, implying $s_{1}(x_{3},x_{1})>s_{3}(x_{3},x_{1})>0$ or $s_{3}(x_{1},x_{3})>s_{1}(x_{1},x_{3})>0$; (iii) $(x_{2},x_{1},x_{3})$, implying $s_{2}(x_{1},x_{2})>s_{1}(x_{1},x_{2})>0$ or $s_{1}(x_{2},x_{1})>s_{2}(x_{2},x_{1})>0$. In each case, $yDx$.
@@
Now suppose $n>3$ and $S\in\mathcal{\widehat{S}}$ is effectively homeo-monotonic. Denote by $(N^{i})_{i=1}^{k}$ and $(A^{i})_{i=1}^{k}$ the respective partitions of $N$ and $A$ into a (possibly empty) top-preference diverse set $N^{1}$ with respect to $A^{1}$ and top-preference-coincident sets $N^{i>1}$ with respect to $A^{i>1}$, $i\leq k$. By the properties of these sets and profile, and by the fourth step of the algorithm, irrespective of the $\gg$-order on $N$ there can be no swaps between $x_{i}$ and $x_{j}$ for $v_{i},v_{j}\in N^{1}$. Thus, $y_{i}=x_{i}$ for all $v_{i}\in N^{1}$. By the properties of $(N^{i>1},A^{i>1})$, moreover and the instruction in Step 4, there can be no swaps between $x_{i}$ and $x_{j}$ for $v_{i}\in N^{l>1}$ and $v_{j}\in N^{m>1}$ when $m,l\in\{2,\ldots,k\}$ are distinct. It follows that $y$ differs from $x$ by at most $n\over 2$} or $n-1\over 2$} pairwise swaps/transpositions if $n$ is even or odd, respectively, with each such $x_{i}\leftrightarrow x_{j}$ swap occurring within the same top-preference-coincident pair $(N^{l>1},A^{l>1})$. By Step 4, finally, every such swap reassigns the—by Pareto-efficiency of xx—jointly preferred item within the relevant pair in the direction of intensity-dominance. Hence, $yDx$. $\square$